# GIM: Improved Interpretability for Large Language Models

### v1

@@ Abstract
Ensuring faithful interpretability in large language models is imperative for trustworthy and reliable AI. A key obstacle is self-repair, a phenomenon where networks compensate for reduced signal in one component by amplifying others, masking the true importance of the ablated component. While prior work attributes self-repair to layer normalization and back-up components that compensate for ablated components, we identify a novel form occurring within the attention mechanism, where softmax redistribution conceals the influence of important attention scores. This leads traditional ablation and gradient-based methods to underestimate the significance of all components contributing to these attention scores. We introduce Gradient Interaction Modifications (GIM), a technique that accounts for self-repair during backpropagation. Extensive experiments across multiple large language models (Gemma 2B/9B, LLAMA 1B/3B/8B, Qwen 1.5B/3B) and diverse tasks demonstrate that GIM significantly improves faithfulness over existing circuit identification and feature attribution methods. Our work is a significant step toward better understanding the inner mechanisms of LLMs, which is crucial for improving them and ensuring their safety. Our code is available at https://github.com/JoakimEdin/gim.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across diverse tasks, from text summarization to code synthesis [1]. Despite their impressive performance, our understanding of their inner mechanisms remains limited. This opacity hinders our ability to explain model outputs, catch hallucinations, systematically improve performance, and ensure reliable behavior in novel situations. To address this, most of the previous explainability methods aim to quantify how internal components (e.g., attention heads, neurons, token embeddings) affect model output [2]. Two of the most widely adopted approaches, perturbation-based and gradient-based methods, both rest on the assumption that modifying a truly salient component will lead to a commensurate change in the model’s output. For instance, perturbation-based methods (e.g., SHAP, LIME, activation patching) investigate the effect of large perturbations [3, 4], while gradient-based methods analyze infinitesimally small perturbations [5]. However, this core assumption that perturbing important components will perturb the model output often fails due to what is known as the self-repair effect [6].
Self-repair is a phenomenon in language models where downstream model components compensate for perturbed components, resulting in deceptively small output changes [7, 8, 6]. As a consequence, explanation methods may underestimate the importance of certain model components [6].
In this paper, we identify a new form of self-repair within the attention mechanism of LLMs. When multiple value vectors associated with large attention weights contain similar information, perturbing any individual attention score has little effect on the model output. If one position’s attention score is reduced, the softmax activation function increases the attention weight of the other positions, maintaining similar output signals (see Figure 1). This attention self-repair systematically causes both perturbation-based methods and gradient methods to underestimate the importance of all components contributing to these attention scores.
To address these challenges, we propose gradient interaction modifications (GIM), a novel gradient-based attribution method specifically designed to account for self-repair effects in language models. GIM introduces targeted modifications to backpropagation to better handle interactions that cause self-repair. First, we develop temperature-adjusted softmax gradients to address the attention self-repair problem. Second, we implement layernorm freeze to deal with the self-repair in layer normalization [9, 6]. Finally, we incorporate gradient normalization as proposed by Achtibat et al. [10], which we found essential for making our backpropagation modifications effective in practice.
@@ 2.1 Self-repair in Neural Networks
Self-repair occurs when ablating components have unexpectedly small effects on model outputs because other components compensate for the ablation. Several studies have identified attention heads that suppress the signal from earlier layers, and the suppression is reduced when those early layers are ablated  [8, 7, 11]. McGrath et al. [8] identified what they called MLP erasure neurons that have a similar suppressing effect on the signal from previous layers. Rushing and Nanda [6] investigated the role of layer normalization in self-repair. They demonstrated that normalization can compensate for ablated layers by rescaling the remaining signals. Our work identifies a new form of self-repair that occurs within attention mechanisms, where softmax redistribution conceals the influence of important attention scores. We also show that this self-repair mechanism can impact the gradient.
@@ 2.2 Gradient-based Attribution Methods
Attribution methods quantify model components’ contribution to specific model behaviors. Feature attribution methods quantify the importance of input features (e.g., tokens), while circuit identification methods quantify the importance of internal model components (e.g., attention heads). Attribution methods are usually perturbation-based or gradient-based [12]. Perturbation-based approaches (e.g., SHAP [3] and activation patching [13]) attribute importance by measuring output changes when perturbing components. However, they require numerous forward passes, making them prohibitively slow for large language models. Gradient-based methods [5, 14] offer a faster alternative by leveraging backpropagation. However, they often produce unfaithful explanations due to neural network non-linearities, where first-order approximations fail to capture true component importance [10, 15]. Methods such as Layer-wise Relevance Propagation (LRP) [16], Integrated Gradients [5], and DeepLIFT [17] attempt to address these limitations by modifying gradient flows, but empirically struggle with language models [18, 19, 10].
Several studies propose specialized backpropagation rules for the transformer architecture. TransformerLRP [9] treats the layer normalization scaling variable and the attention matrix as constants, essentially not backpropagating through them. ATP* [20] propose a backpropagation rule for the softmax operation that better estimates the impact of large attention score perturbations. However, ATP* ignores the normalization contribution in the softmax operation. Achtibat et al. [10] propose AttnLRP, which is a collection of backpropagation rules. Grad norm, their most important contribution to our work, divides the gradient by $N$ when backpropagating through operations where $N$ variables are multiplied together. [^1] Achtibat et al. refer to grad norm as the uniform rule Using Taylor Decomposition and Shapley, they provide theoretical justification for this rule. While TransformerLRP and AttnLRP show that the combination of their backpropagation rules improves the attribution faithfulness, they do not isolate each rule’s contribution to the improvements.
@@ 3 Attention Self-Repair
$\begin{split}&\bm{a}=\bm{Q}_{i}\cdot\bm{K}^{T}\quad\text{(Attention scores for position $i$)}\\ &\bm{s}=\text{Softmax}(\bm{a})\quad\text{(Attention weights for position $i$)}\\ &\bm{o}=\bm{s}\cdot\bm{V}\quad\text{(Attention output for position $i$)}\end{split}$ (1)
@@
Attention mechanisms can be viewed as information routing systems, where each value vector $\bm{V}_{j}$ contain information from position $j$, and the attention weights $\bm{s}_{j}$ determine how much information to copy from that position.
Self-repair occurs when multiple values with large attention weights contain similar information. When an attention score $\bm{a}_{j}$ is ablated, the softmax redistributes weights primarily to positions with the highest remaining scores. If these positions contain values $\bm{V}_{j}$ that contain similar information, the output remains virtually unchanged.
Figure 1 illustrates this with attention weights $[0.5,0.5,0]$ and value vectors $[1,1,0]$. Note that the values are identical at the positions with non-zero attention weights. Individually ablating either attention score (i.e., softmax input) leaves the output unchanged, as the softmax shifts weight to the other position with the same value. Only ablating both scores simultaneously affects the output, as the softmax shifts the attention weights to the position with a different value.
@@ 3.1 Attention self-repair results in zero gradient
To demonstrate the effect of attention self-repair on Equation 3, we first define a set of positions with substantial attention weights $\mathcal{I}_{\epsilon}=\{k:\bm{s}_{k}>\epsilon\}$ where $\epsilon$ is close to zero. Recall that for self-repair to occur, the values $V_{j}$ at the positions with substantial attention weights must contribute similarly to the output. We can measure a value’s unweighted contribution to the output with the dot product between the value and its upstream gradient (GradientXInput). This is equivalent to $\frac{\partial z}{\partial\bm{s}_{j}}$ in Equation 4.
@@
Substituting these into Equation 3, we get:
@@ 4.1 Temperature-adjusted softmax gradients
TSG modifies backpropagation through attention to address the self-repair problem presented in Section 3.1. The goal of TSG is to preserve attribution magnitude when standard backpropagation would otherwise have underestimated the importance of attention scores because of the self-repair effect. TSG addresses the issue with a simple approach:
  1. 1.
    During the forward pass, TSG uses a standard temperature ($T=1$), preserving the model’s normal behavior and predictions
  2. 2.
    During the backward pass, TSG recomputes the softmax with a higher temperature ($T>1$) before computing gradients
Increasing the temperature results in a more uniform attention weight distribution (an increase of entropy). In other words, large attention weights decrease while small weights increase. This increases our set of positions with significant weights $\mathcal{I}_{\epsilon}=\{k:\bm{s}_{k}>\epsilon\}$. Recall that for self-repair to occur, the gradients $\frac{\partial z}{\partial\bm{s}_{j}}$ must be approximately uniform across all positions in $\mathcal{I}_{\epsilon}$. By expanding $\mathcal{I}_{\epsilon}$ to include more positions, we potentially disrupt the uniformity condition that leads to cancellation. When this uniformity condition is broken, the two terms in Equation 3 may no longer completely cancel each other out, allowing for non-zero gradients.
@@ 4.2 Layernorm freeze & gradient norm
Layernorm freeze addresses the self-repair effect in the layer normalization by treating the normalization parameters as constants during backpropagation [6, 9]. Specifically, during the forward pass, we save the normalization factor $\sigma$, and during the backward pass, we divide the upstream gradient by the normalization factor $\sigma$ instead of backpropagating through it. This naively prevents the self-repair effect by preventing layer normalization from changing due to perturbations.
Gradient norm normalizes the gradient after variable multiplications [10]. During backpropagation, it divides the gradient by the number of inputs involved in the multiplication. Achtibat et al. [10] proved the optimality of this approach using Taylor decomposition and Shapley values. Specifically, we divide the gradient by 2 for three key multiplicative interactions in transformer models: 1) attention-value, 2) query-key, and 3) MLP gate-projection.
(a) Difference between ablating the two largest attention scores separately (x-axis) and jointly (y-axis). The 95% quantile interaction is colored orange.
(b) Comparison of gradient and TSG. The attention vectors with the 95% quantile interactions are colored orange.
Figure 2: The attention self-repair effect and how temperature-adjusted softmax gradients increase the gradient when there is self-repair for LLAMA-3.2 1B on FEVER. The left plot compares ablating the two largest attention scores separately and jointly. The strongest interactions are colored orange. The right plot compares the gradient for the same attention scores with the TSG. TSG results in a stronger signal for attention scores with a strong interaction.
@@ 5.1 Models, datasets, and evaluation metrics
For our experiments, we use seven instruction-tuned large language models spanning three model families of various sizes: Gemma-2 (2B, 9B), LLaMA-3.2 (1B, 3B), LLaMA-3.1 (8B), and Qwen-2.5 (1.5B, 3B) [21, 22, 23].
We use six datasets spanning four tasks: BoolQ (question-answering), FEVER and SciFact (fact verification), Twitter sentiment classification and Movie review (sentiment classification), and HateXplain (hate speech detection). These datasets were selected based on their diversity and prevalence in prior attribution studies [24, 25, 12]. We provide more details on the datasets in Section 9.2.
We evaluate the faithfulness of the attributions using Comprehensiveness and Sufficiency [24]. Comprehensiveness ($\uparrow$) measures the average output change after cumulatively ablating top-attributed features:
@@
where $f$ is the model, $\bm{x}$ is the input, $\bm{e}$ contains attribution scores, $p$ is the perturbation function, and $\text{rank}(\cdot)$ orders features by attribution.
Sufficiency ($\downarrow$) measures the average output change after cumulatively ablating the lowest-attributed features:
@@ 5.2 Experiments
1. Attention self-repair in large language models: We investigate whether the attention self-repair problem occurs in large language models and evaluate the effectiveness of TSG in amplifying the signal when gradients would otherwise have underestimated the importance of attention scores.
We first identify attention score vectors where multiple positions contribute significantly to the model output. We define position $j$ as having a significant contribution when:
$\frac{\partial z}{\partial s_{j}}s_{j}>\beta$
where $\beta$ is a threshold parameter.
@@
For each identified attention vector, we measure the joint effect by ablating the two largest attention scores simultaneously versus the sum of their individual ablation effects. If the joint effect is larger than the sum of the individual effect, the attention scores exhibit OR-gate-like behaviour, which is equivalent to self-repair [26]. We quantify these effects by the change in weighted gradient sum $\sum_{k}\frac{\partial z}{\partial s_{k}}s_{k}$. We do this instead of computing the full forward pass to avoid other components compensating for the ablations.
Next, we compare TSG and standard gradient attribution scores for attention scores, focusing on examples with large self-repair effects. This comparison evaluates whether TSG correctly amplifies attribution for attention scores that standard gradients would underestimate due to self-repair.
@@
2. Faithfulness of GIM: For feature attribution evaluation, we compare GIM with five gradient-based methods: GradientXInput (GxI) [5], Integrated gradients (IG) [5], DeepLIFT [27], TransformerLRP [9], and AttnLRP [28]. We did not evaluate perturbation-based methods because they are too computationally expensive for large language models on long inputs [29]. We computed comprehensiveness and sufficiency by replacing tokens with the token representing a space.
@@
For circuit identification evaluation, we compare GIM with attribution patching (ATP) [29], ATP* [^2] We only implemented ATP*’s softmax gradient fix component. [15], and Integrated Gradients [28] by computing comprehensiveness and sufficiency at each layer separately, by replacing the representations by the layer-average representations.
@@
All attribution scores follow the formula $(\bm{x}-\bm{\tilde{x}})\cdot\bm{\alpha}$, where $\bm{x}$ and $\bm{\tilde{x}}$ are original and counterfactual representations, and $\bm{\alpha}$ represents method-specific attributions. For the feature attribution evaluation, we used the token representing a space as the counterfactuals, while for the circuit identification evaluation, we used the layer-average representations as counterfactuals. We used a temperature of 2 for TSG, which we chose based on the results on Gemma-2 2B on the FEVER and HateXplain datasets.
@@
3. Identifying the importance of each GIM modification: Our final experiment is an ablation study, where we estimate the importance of the three modifications in GIM. We estimate their importance by cumulatively adding modifications to standard GradientXInput and measuring the change in feature attribution faithfulness.

### v3

@@ Abstract
Ensuring faithful interpretability in large language models is imperative for trustworthy and reliable AI. A key obstacle is self-repair, a phenomenon where networks compensate for reduced signal in one component by amplifying others, masking the true importance of the ablated component. While prior work attributes self-repair to layer normalization and back-up components that compensate for ablated components, we identify a novel form of self-repair that occurs within the attention mechanism, where softmax redistribution conceals the influence of important attention scores. This attention self-repair leads traditional ablation and gradient-based methods to underestimate the significance of all components contributing to these attention scores. We introduce Gradient Interaction Modifications (GIM), a technique that accounts for self-repair during backpropagation. Extensive experiments across multiple large language models (Gemma 2B/9B, LLAMA 1B/3B/8B, Qwen 1.5B/3B) and diverse tasks demonstrate that GIM significantly improves faithfulness over existing circuit identification and feature attribution methods. Our work is a significant step toward better understanding the inner mechanisms of LLMs, which is crucial for improving them and ensuring their safety. Our code is available at https://github.com/JoakimEdin/gim.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across diverse tasks, from text summarization to code synthesis Brown et al. (2020). Despite their impressive performance, our understanding of their inner mechanisms remains limited. This opacity hinders our ability to explain model outputs, catch hallucinations, systematically improve performance, and ensure reliable behavior in novel situations. To address this, most of the previous explainability methods aim to quantify how internal components (e.g., attention heads, neurons, token embeddings) affect model output Li et al. (2023). Two of the most widely adopted approaches, perturbation-based and gradient-based methods, both rest on the assumption that modifying a truly salient component will lead to a commensurate change in the model’s output. Perturbation-based methods (e.g., SHAP, LIME, activation patching) investigate the effect of large perturbations Lundberg & Lee (2017); Meng et al. , while gradient-based methods analyze infinitesimally small perturbations Sundararajan et al. (2017). However, this core assumption that perturbing important components will perturb the model output often fails due to what is known as the self-repair effect Rushing & Nanda (2024).
Self-repair is a phenomenon in language models where downstream model components compensate for perturbed components, resulting in deceptively small output changes Wang et al. (2022); McGrath et al. (2023); Rushing & Nanda (2024). As a consequence, explanation methods may underestimate the importance of certain model components Rushing & Nanda (2024).
In this paper, we identify a new form of self-repair within the attention mechanism of LLMs. When multiple value vectors associated with large attention weights contain similar information, perturbing any individual attention score has little effect on the model output. If one position’s attention score is reduced, the softmax activation function increases the attention weights of the other positions, maintaining similar output signals (see Figure˜1). This attention self-repair systematically causes both perturbation-based methods and gradient methods to underestimate the importance of all components contributing to these attention scores.
To address these challenges, we propose gradient interaction modifications (GIM), a novel gradient-based attribution method specifically designed to account for self-repair effects in language models. GIM introduces targeted modifications to backpropagation to better handle interactions that cause self-repair. First, we develop temperature-adjusted softmax gradients to address the attention self-repair problem. Second, we implement layernorm freeze to deal with the self-repair in layer normalization Ali et al. (2022); Rushing & Nanda (2024). Finally, we incorporate gradient normalization as proposed by Achtibat et al. (2024), which we found essential for making our backpropagation modifications effective in practice.
@@ 2.1 Self-repair in Neural Networks
Self-repair occurs when ablating components has unexpectedly small effects on model outputs because other components compensate for the ablation. Several studies have identified attention heads that suppress the signal from earlier layers, and the suppression is reduced when those early layers are ablated  McGrath et al. (2023); Wang et al. (2022); McDougall et al. (2023). McGrath et al. (2023) identified what they called MLP erasure neurons that have a similar suppressing effect on the signal from previous layers. Rushing & Nanda (2024) investigated the role of layer normalization in self-repair. They demonstrated that normalization can compensate for ablated layers by rescaling the remaining signals. Our work identifies a new form of self-repair that occurs within attention mechanisms, where softmax redistribution conceals the influence of important attention scores. We also show that this self-repair mechanism can impact the gradient.
@@ 2.2 Gradient-based Attribution Methods
Attribution methods quantify model components’ contribution to specific model behaviors. Feature attribution methods quantify the importance of input features (e.g., tokens), while circuit identification methods quantify the importance of internal model components (e.g., attention heads). Most attribution methods are perturbation-based or gradient-based Lyu et al. (2024). Perturbation-based approaches (e.g., SHAP Lundberg & Lee (2017) and activation patching Meng et al. (2022)) attribute importance by measuring output changes when perturbing components. However, they require numerous forward passes, making them prohibitively slow for large language models. Gradient-based methods offer a faster alternative by leveraging backpropagation Sundararajan et al. (2017); Simonyan et al. (2014). However, they often produce unfaithful explanations due to neural network non-linearities, where first-order approximations fail to capture true component importance Achtibat et al. (2024); Kramár et al. (2024a). Methods such as Layer-wise Relevance Propagation (LRP) Bach et al. (2015), Integrated Gradients Sundararajan et al. (2017), and DeepLIFT Shrikumar et al. (2017b) attempt to address these limitations by modifying gradient flows, but empirically struggle with language models Edin et al. (2025, 2024); Achtibat et al. (2024).
Several studies propose specialized backpropagation rules for the transformer architecture. TransformerLRP Ali et al. (2022) treats the layer normalization scaling variable and the attention matrix as constants, essentially not backpropagating through them. ATP* Kramár et al. (2024b) propose a backpropagation rule for the softmax operation that better estimates the impact of large attention score perturbations. However, ATP* ignores the normalization contribution in the softmax operation. Achtibat et al. (2024) propose AttnLRP, which is a collection of backpropagation rules. Grad norm, their most important contribution to our work, divides the gradient by $N$ when backpropagating through operations where $N$ variables are multiplied together. [^1] Achtibat et al. refer to grad norm as the uniform rule Using Taylor Decomposition and Shapley, they provide theoretical justification for this rule. While TransformerLRP and AttnLRP show that the combination of their backpropagation rules improves the attribution faithfulness, they do not isolate each rule’s contribution to the improvements.
@@ 3 Attention Self-Repair
[…]
@@
Attention mechanisms can be viewed as information routing systems, where each value vector $\bm{V}_{j}$ contains information from position $j$, and the attention weights $\bm{s}_{j}$ determine how much information to copy from that position.
Self-repair occurs when multiple values with large attention weights contain similar information. When an attention score $\bm{a}_{j}$ is ablated, the softmax redistributes weights primarily to positions with the highest remaining scores. If these positions contain values $\bm{V}_{j}$ that contribute similar information, the output remains virtually unchanged.
Figure˜1 illustrates this with attention weights $[0.5,0.5,0]$ and value vectors $[1,1,0]$. Note that the values are identical at the positions with non-zero attention weights. Individually ablating either attention score (i.e., softmax input) leaves the output unchanged, as the softmax shifts the weight to the other position with the same value. Only ablating both scores simultaneously affects the output, as the softmax shifts the attention weights to the position with a different value.
@@ 3.1 Attention self-repair results in zero gradient
To demonstrate the effect of attention self-repair on Equation˜3, we first define a set of positions with substantial attention weights $\mathcal{I}_{\epsilon}=\{k:\bm{s}_{k}>\epsilon\}$ where $\epsilon$ is close to zero. Recall that for self-repair to occur, the values $V_{j}$ at the positions with substantial attention weights must contribute similarly to the output. We can measure a value’s unweighted contribution to the output with the dot product between the value and its upstream gradient (GradientXInput). This is equivalent to $\frac{\partial z}{\partial\bm{s}_{j}}$ in Equation˜4.
@@
Substituting these into Equation˜3, we get:
@@ 4.1 Temperature-adjusted softmax gradients
TSG modifies backpropagation through attention to address the self-repair problem presented in Section˜3.1. The motivation behind TSG is well explained through the "firing squad" analogy from Pearl (2009): two soldiers fire at a prisoner, and we want to measure the causal importance of each soldier for the prisoner’s death. If we measure this by separately preventing each soldier from firing, we would incorrectly conclude that both are innocent, since either shot alone is fatal. To measure their true responsibility, we must simultaneously prevent both shooters from firing.
Attention self-repair exhibits similar OR-gate behavior to the firing squad example. The gradient of the output with respect to the attention scores measures the effect of perturbing each attention score separately. Instead, we designed TSG to approximate the effect of perturbing multiple attention scores simultaneously. We developed TSG empirically by evaluating which techniques led to the fastest and most accurate approximation of jointly ablating multiple attention scores, while also producing faithful explanations. We will later demonstrate the accuracy of this approximation.
During backpropagation, TSG recomputes the softmax with a higher temperature ($T>1$) before computing the softmax gradients. This results in a more uniform distribution of attention weights: large attention weights decrease while small weights increase. This increases our set of positions with significant weights $\mathcal{I}_{\epsilon}=\{k:\bm{s}_{k}>\epsilon\}$. Recall that for self-repair to occur, the gradients $\frac{\partial z}{\partial\bm{s}_{j}}$ must be approximately uniform across all positions in $\mathcal{I}_{\epsilon}$. By expanding $\mathcal{I}_{\epsilon}$ to include more positions, we increase the probability of a value vector disrupting the uniformity condition. When this uniformity condition is broken, the two terms in Equation˜3 no longer completely cancel each other out, allowing for non-zero gradients.
@@ 4.2 Layernorm freeze & gradient norm
Layernorm freeze addresses the self-repair effect in the layer normalization by treating the normalization parameters as constants during backpropagation Rushing & Nanda (2024). Specifically, during the forward pass, we save the normalization factor $\sigma$, and during the backward pass, we divide the upstream gradient by the normalization factor $\sigma$ instead of backpropagating through it. This naively prevents the self-repair effect by preventing layer normalization from changing due to perturbations.
Gradient norm normalizes the gradient after variable multiplications Achtibat et al. (2024). During backpropagation, it divides the gradient by the number of inputs involved in the multiplication. Achtibat et al. (2024) proved the optimality of this approach using Taylor decomposition and Shapley values. Specifically, we divide the gradient by 2 for three key multiplicative interactions in transformer models: 1) attention-value, 2) query-key, and 3) MLP gate-projection.
@@ 5 Experimental setup
Figure 2: The attention self-repair effect and how temperature-adjusted softmax gradients approximate the joint ablation effect for LLAMA-3.2 1B on FEVER. The figures compare ablating the two largest attention scores jointly with a) ablating them separately, b) gradients, and c) TSG.
@@ 5.1 Models, datasets, and evaluation metrics
For our experiments, we use seven instruction-tuned large language models spanning three model families of various sizes: Gemma-2 (2B, 9B), LLaMA-3.2 (1B, 3B), LLaMA-3.1 (8B), and Qwen-2.5 (1.5B, 3B) Team et al. (2024); Grattafiori et al. (2024); Qwen et al. (2025).
We use six datasets spanning four tasks: BoolQ (question-answering), FEVER and SciFact (fact verification), Twitter sentiment classification and Movie review (sentiment classification), and HateXplain (hate speech detection). These datasets were selected based on their diversity and prevalence in prior attribution studies Lyu et al. (2024). We provide more details on the datasets in Section˜A.3.
We evaluate the faithfulness of the attributions using Comprehensiveness and Sufficiency DeYoung et al. (2020).
Comprehensiveness ($\uparrow$ the higher the better) measures the average output change after cumulatively ablating top-attributed features:
@@
where $f$ is the model, $\bm{x}$ is the input, $\bm{e}$ contains attribution scores, $N$ is the number of input features, $p$ is the perturbation function, and $\text{rank}(\cdot)$ orders features by attribution.
Sufficiency ($\downarrow$ the lower the better) measures the average output change after cumulatively ablating the lowest-attributed features:
@@ 5.2 Experiments
Attention self-repair in large language models: We investigate whether the attention self-repair problem occurs in large language models and evaluate the effectiveness of TSG in approximating the effect of perturbing multiple attention scores simultaneously.
To identify instances of self-repair in practice, we apply the following procedure. We focus on the top $1\%$ most important attention weights, which we compute as $\frac{\partial z}{\partial s_{j}}s_{j}$. From these, we select only attention weight vectors that have multiple significant attention weights (each >0.01>0.01). Finally, we classify these cases as exhibiting self-repair if the value vectors at positions with significant attention weights are similar to each other. To identify similar value vectors, we first compute their contributions to the model output: $\frac{\partial z}{\partial o}V_{j}$. We then calculate the coefficient of variation (standard deviation divided by mean) across these contributions. Intuitively, a low coefficient of variation indicates that the value vectors contribute similarly to the output, which is the condition under which self-repair occurs, and gradients cancel out. If this value is less than $0.1$, we classify the case as self-repair.
@@
For each identified self-repair in attention, we measure the joint effect by ablating the two largest attention scores simultaneously versus the sum of their individual ablation effects. If the joint effect is larger than the sum of the individual effects, the attention scores exhibit OR-gate-like behaviour, which is equivalent to self-repair Tsang et al. (2020). We quantify these effects by the change in weighted gradient sum $\sum_{k}\frac{\partial z}{\partial s_{k}}s_{k}$. We do this instead of computing the full forward pass to avoid other components compensating for the ablations. Finally, we show that TSG better approximates the joint effect than standard gradients by comparing both with the joint effects.
@@
Faithfulness of GIM: For feature attribution evaluation, we compare GIM with five gradient-based methods: GradientXInput (GxI) Sundararajan et al. (2017), Integrated gradients (IG) Sundararajan et al. (2017), DeepLIFT Shrikumar et al. (2017a), TransformerLRP Ali et al. (2022), and AttnLRP Hanna et al. (2024). We did not evaluate perturbation-based methods because they are too computationally expensive for large language models on long inputs Syed et al. (2023). We computed comprehensiveness and sufficiency by replacing tokens with the token representing a space.
@@
For circuit identification evaluation, we compare GIM with attribution patching (ATP) Syed et al. (2023), ATP* [^2] We only implemented ATP*’s softmax gradient fix component. Kramár et al. (2024a), and Integrated Gradients Hanna et al. (2024) by computing comprehensiveness and sufficiency at each layer separately, by replacing the representations with the layer-average representations. We also evaluate GIM on the Mechanistic Interpretability Benchmarks Mueller et al. (2025). We provide more details in Section˜A.1.
@@
All attribution scores follow the formula $(\bm{x}-\bm{\tilde{x}})\cdot\bm{\alpha}$, where $\bm{x}$ and $\bm{\tilde{x}}$ are original and counterfactual representations, and $\bm{\alpha}$ represents method-specific attributions. For the feature attribution evaluation, we used the token representing a space as the counterfactuals. For the circuit identification evaluation, we used the layer-average representations as counterfactuals. We used a temperature of 2 for TSG, which we chose based on the results on Gemma-2 2B on the FEVER and HateXplain datasets.
@@
Identifying the importance of each GIM modification: Our final experiment is an ablation study, where we estimate the importance of the three modifications in GIM. We measure the change of feature attribution faithfulness when cumulatively adding modifications to the standard GradientXInput.